# ProtoSiTex: Learning Semi-Interpretable Prototypes for Multi-label Text Classification

### v1

@@ 4.2. Evaluation Metrics
Figure 3. Qualitative analysis of ProtoSiTex on samples from (a) HR, (b) TweetEVAL (Barbieri et al., 2020), and (c) IMDb (Maas et al., 2011) datasets
4.3. Results & Comparative Study
To compare with ProtoSiTex, we chose 6 recent prototypical learning models: ProSeNet (Ming et al., 2019), ProtoTEX (Das et al., 2022), ProtoryNet (Hong et al., 2023), Proto-lm (Xie et al., 2023), GAProtoNet (Wen et al., 2025), and ProtoLens (Wei et al., 2025). In addition to this, we selected 9 black-box LM performances: ALBERT (Lan et al., 2020), DeBERTaV3 (He et al., 2023) BERT (Devlin et al., 2019), DistilBERT (Sanh et al., 2019), ELECTRA (Clark et al., 2020), XLNET (Yang et al., 2019), ModernBERT (Warner et al., 2025) RoBERTa and RoBERTa-Large (Liu et al., 2019) and 2 text classification methods: TextGCN (Yao et al., 2019b) and BertGCN (Lin et al., 2021), without incorporating interpretability.
In Table 2, we present the comparative study. This study reveals that existing prototype-based SOTA approaches often sacrifice performance for interpretability, leading to significantly lower results compared to baselines, particularly on challenging multi-label datasets such as HR. In contrast, ProtoSiTex successfully overcame this limitation by narrowing the performance gap with black-box LMs, while preserving inherent interpretability across varying dataset complexities. Specifically, ProtoSiTex achieved improvements of 0.04% and 2.85% in $\mathcal{A}$} on the IMDb and TweetEVAL datasets, respectively, and delivered a substantial 12.93% gain in $\mathcal{F}_{m}$} on HR, compared to the best prior prototype-based methods. These results across three diverse tasks highlight ProtoSiTex’s generalizability and applicability.
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$75758080858590909595K-MeansKaiming NormXavier NormRandom Norm
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$6060707080809090100100w/o ATProtoSiTex
$\mathcal{F}$$\mathcal{BA}$${\mathcal{HL}}^{c}$2020404060608080100100Euclidean Dist.Cosine Sim.ProtoSiTex
[…]
Figure 4. Performance analysis of ProtoSiTex on HR dataset: (a) with various prototype initializers; (b) without alternate training (w/o AT); (c) across different similarity measures; impact of the number of (d) prototypes, and (e) heads
[…]
Figure 5. Impact of hyperparameters $\alpha_{i}$ and $\lambda_{i}$
4.4. Qualitative Analysis of ProtoSiTex
To illustrate ProtoSiTex’s interpretability and prediction quality, we present qualitative examples from the HR, TweetEVAL, and IMDb datasets (Fig. 3). In each case, the model accurately predicts aspect or sentiment labels while grounding them in semantically coherent sub-sentence segments via learned prototypes. In the HR example, ProtoSiTex correctly identifies multi-label aspects (Room & Amenities, Food, Pricing & Value, etc.) from a single review, while providing text-grounded justifications. This highlights its ability to disentangle overlapping review aspects and attribute them to distinct prototype-based explanations. In the TweetEVAL sample, ProtoSiTex captures the emotion joy and connects it to meaningful prototypes reflecting positive sentiment, demonstrating its effectiveness in emotional reasoning with minimal text context. Similarly, in the IMDb example, ProtoSiTex predicts the positive label and supports it with a rich prototype cluster centered on artistic and narrative appreciation, confirming alignment with human sentiment. Overall, these examples demonstrate that ProtoSiTex offers not only accurate predictions but also transparent, faithful interpretations at the sub-sentence level, making it suitable for real-world applications requiring interpretability. More qualitative results are shown in Appendix B.
4.5. Impact of Model Components
Due to space constraints, we report the impact of the module only on the HR dataset, which is representative enough to illustrate the observed trends.
4.5.1. Impact of Different Prototype Initializers:
We evaluated four prototype initialization strategies: random, K-Means centroids, Kaiming Normal, and Xavier Normal. As shown in Fig. 4(a), random initialization consistently outperformed the others across $\mathcal{F}_{m}$, $\mathcal{BA}_{m}$, and ${\mathcal{HL}}^{c}$. This performance gain is attributed to the increased diversity introduced early in the training, which helps avoid premature convergence and overfitting.
4.5.2. Impact of Alternate Training:
ProtoSiTex adopts a dual-phase training strategy, alternating between unsupervised prototype learning and supervised classification. To evaluate its efficacy, we compared model variants trained with and without alternation. As shown in Fig. 4(b), alternating training significantly improved performance by enabling the model to discover semantically meaningful prototype clusters and align them with downstream classification objectives. In contrast, without alternation, the model struggled to learn task-specific prototypes, lowering performance.
4.5.3. Impact of Similarity Function:
4.6. Impact of Hyperparameters
ProtoSiTex includes several key hyperparameters: the number of prototypes ($q$), number of heads ($h$), loss weights for clustering ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$), and classification ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$). To systematically investigate the influence of these hyperparameters on model performance, we conducted a series of experiments on the HR dataset. As shown in Fig. 4(d), increasing $q$ initially boosts performance by capturing diverse semantic concepts, but excessive values lead to redundancy and over-smoothing. A similar trend is observed for $h$ in Fig. 4(e). Furthermore, Fig. 5 shows that assigning $\alpha_{1}=0.01$, $\alpha_{2}=0.98$, $\alpha_{3}=0.01$, $\lambda_{1}=0.6$, $\lambda_{2}=0.2$, and $\lambda_{3}=0.2$ yields optimal results. This multi-level supervision prevented the model from overfitting to a single level of granularity, supporting both local and global interpretability and accuracy.
[…]
Figure 6. Aspect-wise performance of ProtoSiTex on HR
@@ 4.6. Impact of Hyperparameters
Figure 7. Misprediction examples of ProtoSiTex on (a) HR, (b) TweetEVAL (Barbieri et al., 2020), and (c) IMDb (Maas et al., 2011) datasets
4.7. Aspect-wise Performance Analysis:
Fig. 6 presents a detailed aspect-wise evaluation of ProtoSiTex on the HR dataset, offering insights into the model’s performance across individual aspect categories. Notably, the aspects Guest Experience, Food, and Concierge not only appear the most frequently in HR but also exhibit strong predictive performance. In contrast, Pricing & Value and Personalization are the least represented. Despite the challenges posed by class imbalance, ProtoSiTex shows a strong ability to identify the Pricing & Value aspect, achieving $\mathcal{F}$ and $\mathcal{A}$ scores comparable to those of the more frequent categories. However, Personalization remains the most challenging aspect, particularly in terms of $\mathcal{R}$, as the model often fails to capture all relevant instances. This difficulty is likely due to both its low frequency in the dataset and its inherently subjective nature. Unlike more explicit aspects such as Food or Location & Surrounding, Personalization is often expressed through subtle, indirect, and context-dependent language, making it more difficult to detect.
4.8. Misprediction Analysis
@@
(i) Boundary Ambiguity in Prototype: In multi-label classification, ProtoSiTex occasionally exhibits boundary ambiguity among prototypes, where semantically related class representations capture overlapping evidence. This ambiguity leads to confusion in distinguishing co-occurring categories, resulting in the assignment of only a single dominant label instead of multiple relevant ones. For example, in Fig. 7(a), the review “Spacious room with balcony with amazing view of the valley and the fall”. The actual labels are Room & Amenities and Location & Surrounding. However, ProtoSiTex predicts only Room & Amenities. Here, Room & Amenities dominates and captures shared contextual characteristics such as “spacious room”, “balcony”, overshadowing the equally valid cues of “view of the valley” that should activate Location & Surrounding.
@@
(ii) Lexical cue dominance: Sometimes the mispredictions in ProtoSiTex were caused by lexical bias, the tendency to rely excessively on high-intensity words when making predictions. Words such as “angry”, “worst”, or “best” often dominate the model’s decision process, even when their contextual meaning is softened/inverted by surrounding text. For instance, in Fig. 7(b) the review “i love ur angry comments”, the actual emotion is affectionate teasing or joy, yet the ProtoSiTex incorrectly predicts anger due to the presence of the word “angry”.
@@
(iii) Absence of Fine-Grained Annotation: This observation highlights an important takeaway: even when only coarse document-level supervision is available, prototype-based explanations remain highly useful for error analysis and model trust. By showing how individual sub-sentences align with prototypes tied to different target classes, the model offers interpretable evidence that pinpoints why a misclassification occurred, rather than treating it as a black-box error. From a technical perspective, the misclassification arises due to aggregation bias, while sub-sentence–prototype mappings are accurate, the hierarchical aggregation process (from prototype-level activations to the final document-level decision) can overweight strong local cues and misjudge the overall label. In practice, this means that even when the document-level prediction is wrong, the prototype-level reasoning still faithfully explains the decision path, helping to uncover issues such as prototype imbalance, lack of discourse-level cues (e.g., concessions or sarcasm), or dataset limitations tied to coarse supervision. Thus, beyond accuracy metrics, our approach demonstrates that interpretable prototype assignments provide actionable insights for debugging and model refinement. As illustrated in Fig. 7(c), the identified prototypes, explanations, and associated class mappings are correct, even though the final document-level prediction was misclassified.
Conclusion
@@
We proposed ProtoSiTex, a semi-interpretable framework for fine-grained, multi-label text classification. By combining adaptive prototype learning, dual-phase alternating training, and hierarchical supervision, ProtoSiTex delivers accurate and transparent predictions across multiple levels of granularity. To support evaluation, we introduced the Hotel Reviews (HR) dataset with sub-sentence-level multi-aspect annotations. Extensive experiments on HR, along with the publicly available IMDb and TweetEVAL datasets, show that ProtoSiTex matches or exceeds baseline and state-of-the-art models, while providing faithful, human-aligned interpretations. Future directions include extending ProtoSiTex to cross-domain and multilingual settings, incorporating multimodal cues, and integrating large language models for interactive, user-centric explanations.
Appendix A Hyperparameters Configuration
Table 3. Brief details of hyperparameters
Attribute Value
Optimizer AdamW
Learning rate (Clustering-phase) $2\times 10^{-5}$
Learning rate (Classification-phase) $2\times 10^{-5}$
Batch size 64
LM encoder (RoBERTa-Large): Feature dimension 1024
No. of dense layer 2
Dimension of dense layer 256 & 64
Clustering loss parameters ($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ ) [0-1]
Classification loss parameters ($\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$) [0-1]
No. of head [1-32]
No. of prototype [4-64]
Table 3 presents the hyperparameters engaged in ProtoSiTex. To obtain high-dimensional, contextually rich text embeddings, we incorporate a language model, called RoBERTa (Liu et al., 2019). Specifically, it is a robustly optimized BERT variant, which we employed in its large-weight version pretrained on large English-language internet corpora using the masked language modeling objective. RoBERTa-large remains tunable within our framework. We maintain a consistent dropout rate of 0.1 across all trainable weights. To avoid overfitting, early stopping with a patience of three epochs based on validation performance is applied. Additionally, we set a maximum training limit of 100 epochs. Each epoch includes 10 iterations of the clustering phase and 100 iterations of the classification phase, ensuring a balanced optimization of both phases.
Appendix B Qualitative Analysis
While we presented a qualitative analysis in the main paper, this section provides a more detailed examination, offering finer-grained insights into the interpretability and effectiveness of ProtoSiTex. Tables 4 show textual representation of learned prototypes along with their corresponding classes, derived after training on all datasets (Hotel Reviews (HR), IMDb (Maas et al., 2011), and TweetEVAL (Barbieri et al., 2020)). Our analysis reveals that the learned prototypes are well-distributed across each dataset and span all labels. Additionally, we show qualitative results for samples from HR, IMDb, and TweetEVAL in Fig.s 8.
B.1. Analysis on Hotel Reviews (HR)
Fig. 8(a) demonstrates the interpretability of ProtoSiTex on multi-aspect text. Here, the hotel review referencing honeymoon context directly supports the prototype of “Fostering Memorable Celebrations and Immersive Activities”, while expressions such as “hotel has the best view to offer” closely match the prototype “Location: Scenic and Strategically Advantageous Settings”. ProtoSiTex also provides rich prototype justification for additional aspects and successfully predicts the review into Concierge, Personalization, Food, Location & Surrounding, Guest Experience aspects.
B.2. Analysis on IMDb
In Fig. 8(b), ProtoSiTex classifies the input text as positive, recognizing its emphasis on how this film version outshines all previous adaptations. The model captures the sentiment of novelty and strength in storytelling, with a strong alignment to relevant prototypes, demonstrating both the interpretability and effectiveness of ProtoSiTex. Similarly, ProtoSiTex predicts negative for the sample presented in Fig. 8(c).
B.3. Analysis on TweetEval
Fig. 8(d-g) illustrate the two-way interpretability of ProtoSiTex on TweetEVAL. Specifically, Fig. 8(d), the input text reflects deep emotional pain rooted in past trauma, conveying feelings of depression and grief. This aligns closely with the prototype capturing themes of “Existential dread and loss”, showcasing ProtoSiTex’s ability to contextualize emotional content through semantically coherent prototypes.
(a)Refer to caption
(b) Refer to caption (c) Refer to caption
(d) Refer to caption (e) Refer to caption
(f) Refer to caption (g) Refer to caption
Figure 8. ProtoSiTex qualitative analysis on samples from (a) HR, (b)-(c) IMDb (Maas et al., 2011), and (d)-(g) TweetEVAL (Barbieri et al., 2020) datasets
Table 4. Textual representative of learned prototypes on All Datasets
Sl. Prototype Texts Labels HR Dataset (Ours) 1 Arrival logistics, personalization, and service recovery Concierge, Personalization, Guest Experience 2 Personalization and navigating external factors Personalization, Guest Experience 3 Inconsistent functionality and safety concerns Room & Amenities, Guest Experience 4 Complimentary room upgrades and enhanced stay benefits Personalization, Room & Amenities 5 Exceptional guest experience through proactive service and unexpected delights Guest Experience, Concierge 6 Room quality and prime location versus maintenance and staff issues Room & Amenities, Location & Surrounding, Concierge 7 Housekeeping: excellence in service marred by pest issues Housekeeping 8 Room features, accessibility, and guest privacy concerns. Room & Amenities, Guest Experience 9 Varied impressions of hotel property and service Guest Experience, Room & Amenities 10 State of facilities and guest offerings Room & Amenities 11 Assessing cost-effectiveness and perceived worth Pricing & Value 12 Well-maintained facilities and friendly service. Room & Amenities, Concierge 13 Exceptional staff and room quality Room & Amenities, Housekeeping 14 Location: proximity, environmental noise, and scenic aspects Location & Surrounding 15 Navigating operational shortcomings and external challenges Guest Experience, Concierge 16 Cleanliness and quality: rooms, amenities, and breakfast Housekeeping, Room & Amenities, Food 17 Inconsistent guest experience and room and amenities quality Room & Amenities, Guest Experience 18 Varied staff performance and property conditions with notable location and dining Guest Experience, Room & Amenities, Food, Location & Surrounding 19 Variable quality across guest experience and property aspects Guest Experience, Room & Amenities 20 Balancing location advantages with pricing concerns Location & Surrounding, Pricing & Value 21 Overpriced with inadequate service and amenities Pricing & Value, Room & Amenities 22 Recurring room and amenity malfunctions and service deficiencies Room & Amenities, Housekeeping 23 Financial dealings and perceived value Pricing & Value 24 Inconsistent value for money with unexpected additional costs Pricing & Value 25 Spacious and family-friendly Room & Amenities 26 Quality and utility of facilities and in-room provisions Room & Amenities 27 Varied room types, upgrades, and amenity performance Room & Amenities 28 Diverse guest insights across hotel operations, location ambiance, and service delivery Guest Experience, Location & Surrounding, Concierge 29 Location: scenic and strategically advantageous settings Location & Surrounding 30 Contrasting guest experience and perceived value Guest Experience, Pricing & Value 31 Service and billing management Concierge, Pricing & Value 32 Elevated stays through thoughtful amenities and diverse culinary offerings Room & Amenities, Food, Guest Experience 33 Discrepant billing, payment obstacles, and perceived poor value Pricing & Value 34 Outstanding hospitality, culinary delights, and comfort. Food, Concierge 35 Varying quality in food, facilities, and service Food 36 Focus on renovation and comfort upgrades Room & Amenities 37 Handling of booking, billing, and diverse guest requests Concierge 38 Property condition, aesthetic appeal, and service quality Room & Amenities 39 Prominent on-site dining and accommodation features Food, Room & Amenities 40 Diverse dining experiences and efficient guest services Food , Concierge 41 Fostering memorable celebrations and immersive activities Guest Experience 42 Mixed perceptions of room condition, amenities, and service quality relative to cost Room & Amenities, Guest Experience, Pricing & Value 43 Location: a decisive factor for guests Location & Surrounding 44 Hotel’s location and on-site environment: accessibility, amenities, and challenges Location & Surrounding, Room & Amenities 45 Food & beverage: value, variety, and pricing transparency Food, Pricing & Value 46 Staff engagement and support Concierge 47 Personalized attention and special arrangements Personalization, Concierge 48 Modern amenities and varied room conditions with notable maintenance and service lapses Room & Amenities, Housekeeping, Guest Experience IMDB Dataset (Maas et al., 2011) 1 Unveiling the depth of human emotion through diverse narratives Positive 2 Discerning appraisals of cinematic,theatrical performances, adaptations Positive 3 Celebrating unconventional and engrossing storytelling Positive 4 Potent narratives of human struggle and unflinching reality Positive 5 Diverse perspectives and defended convictions Positive 6 Cinema of intense debate and baffling choices Positive 7 Memorable, albeit flawed, cinematic experiences Negative 8 Films with commendable elements, yet falling short of their potential Negative TweetEVAL Dataset (Barbieri et al., 2020) 1 Irritation and frustration Anger 2 Anger and disgust Anger 3 Irritation and disrespect triggered outbursts Anger 4 Humorous and offended outbursts Anger 5 Contentment and reflection Joy 6 Cuddling and positivity promote well-being Joy 7 Gratitude and joyful appreciation Joy 8 Regret, resilience, and hope amidst adversity Optimism 9 Minor discomforts and empathetic support Optimism 10 Gloomy discomfort Sadness 11 Disappointment and discouragement Sadness 12 Existential dread and loss Sadness

### v2

@@ IV-F Impact of Hyperparameters
Figure 6: Correctly predicted examples of ProtoSiTex on (a) HR, (b) TweetEVAL [TweetEVAL], and (c) IMDb [IMDb]. Best viewed in color.
@@
Figure 7: Misprediction examples of ProtoSiTex on (a) HR, (b) TweetEVAL [TweetEVAL], and (c) IMDb [IMDb]. Best viewed in color.
IV-G Aspect-wise Performance Analysis
Fig. 5 presents a detailed aspect-wise evaluation of ProtoSiTex on the HR dataset, offering insights into the model’s performance across individual aspect categories. Notably, the aspects Guest Experience, Food, and Concierge not only appear the most frequently in HR but also exhibit strong predictive performance. In contrast, Pricing & Value and Personalization are the least represented. Despite the challenges posed by class imbalance, ProtoSiTex shows a strong ability to identify the Pricing & Value aspect, achieving $\mathcal{F}$ and $\mathcal{A}$ scores comparable to those of the more frequent categories. However, Personalization remains the most challenging aspect, particularly in terms of $\mathcal{R}$, as the model often fails to capture all relevant instances. This difficulty is likely due to both its low frequency in the dataset and its inherent subjective nature. Unlike more explicit aspects such as Food or Location & Surrounding, Personalization is often expressed through subtle, indirect, and context-dependent language, making it more difficult to detect.
IV-H Qualitative Analysis of ProtoSiTex
To illustrate ProtoSiTex’s interpretability and prediction quality, we present correctly predicted examples from the HR, TweetEVAL, and IMDb datasets in Fig. 6. In each case here, the model accurately predicts aspect or sentiment labels while grounding them in semantically coherent subsentence segments via learned prototypes. In the HR example, ProtoSiTex correctly identifies multi-label aspects (Room & Amenities, Food, Pricing & Value, etc.) from a single review, while providing text-grounded justifications. This highlights its ability to disentangle overlapping review aspects and attribute them to distinct prototype-based explanations. In the TweetEVAL sample, ProtoSiTex captures the emotion joy and connects it to meaningful prototypes reflecting positive sentiment, demonstrating its effectiveness in emotional reasoning with minimal text context. Similarly, in the IMDb example, ProtoSiTex predicts the positive label and supports it with a rich prototype cluster centered on artistic and narrative appreciation, confirming alignment with human sentiment. Overall, these examples demonstrate that ProtoSiTex offers not only accurate predictions but also transparent, faithful interpretations at the subsentence level, making it suitable for real-world applications requiring interpretability.
IV-I Misprediction Analysis
@@ IV-I1 Boundary Ambiguity in Prototype
In multi-label classification, ProtoSiTex occasionally exhibits boundary ambiguity among prototypes, where semantically related class representations capture overlapping evidence. This ambiguity leads to confusion in distinguishing co-occurring categories, resulting in the assignment of only a single dominant label instead of multiple relevant ones. For example, in Fig. 7(a), the review “Spacious room with balcony with amazing view of the valley and the falls”. The actual labels are Room & Amenities and Location & Surrounding. However, ProtoSiTex predicts only Room & Amenities. Here, Room & Amenities dominates and captures shared contextual characteristics such as “spacious room”, “balcony”, overshadowing the equally valid cues of “view of the valley and the falls” that should activate Location & Surrounding.
IV-I2 Lexical Cue Dominance
@@
Sometimes the mispredictions in ProtoSiTex were caused by lexical bias, the tendency to rely excessively on high-intensity words when making predictions. Words such as “angry”, “worst”, or “best” often dominate the model’s decision process, even when their contextual meaning is softened/ inverted by surrounding text. For instance, in Fig. 7(b) the review “@user i love ur angry comments”, the actual emotion is affectionate teasing or joy, yet the ProtoSiTex incorrectly predicts anger due to the presence of the word “angry”.
IV-I3 Absence of Fine-Grained Annotation
@@
This observation highlights an important takeaway: even when only coarse document-level supervision is available, prototype-based explanations remain highly useful for error analysis and model trust. By showing how individual subsentences align with prototypes tied to different target classes, ProtoSiTex offers interpretable evidence that pinpoints why a misclassification occurred, rather than treating it as a black-box error. From a technical perspective, the misclassification arises due to aggregation bias, while subsentence–prototype mappings are accurate, the hierarchical aggregation process (from prototype-level activations to the final document-level decision) can overweight strong local cues and misjudge the overall label. In practice, this means even when the document-level prediction is wrong, the prototype-level reasoning still faithfully explains the decision path, helping to uncover issues such as prototype imbalance, lack of discourse-level cues, or dataset limitations tied to coarse supervision. Thus, beyond accuracy metrics, our approach demonstrates that interpretable prototype assignments provide actionable insights for debugging and model refinement. As illustrated in Fig. 7(c), the identified prototypes, explanations, and associated class mappings are correct, even though the final document-level prediction was misclassified.
Additional experimental analyses are provided in the supplementary file. Appendix C presents the convergence behavior of the individual and joint loss functions during training. Appendix D offers a comprehensive evaluation of prototype interpretability using multiple metrics, including coverage, contrastivity, centered kernel alignment, and fidelity. Appendix E reports a detailed nonparametric statistical significance analysis based on the Friedman and Nemenyi tests. Finally, Appendix F provides extensive qualitative results, illustrating representative prototypes generated across the three benchmark datasets.
V Conclusion
@@
We proposed ProtoSiTex, a semi-interpretable framework for fine-grained, multi-label text classification. By combining adaptive prototype learning, dual-phase alternate training, and hierarchical supervision, ProtoSiTex delivers accurate and transparent predictions across multiple levels of granularity. To support evaluation, we introduced the hotel reviews dataset (HR) with subsentence-level multi-aspect annotations. Extensive experiments on HR, along with the publicly available IMDb and TweetEVAL datasets, show that ProtoSiTex matches or exceeds baseline and state-of-the-art models, while providing faithful, human-aligned interpretations. Future directions include extending ProtoSiTex to cross-domain and multilingual settings, incorporating multimodal cues, and integrating large language models for interactive, user-centric explanations.
Acknowledgment
The authors heartily thank all the researchers/ interns/ annotators who helped in hotel reviews dataset preparation.